### v4

@@ 3.3 Customized Distribution Change Surrogate
[…]
@@ 5.1 Performance Analysis among Representative Methods
[…]
@@ 6 Related Works
Simulation of Distribution Changes in Machine Learning. Distribution changes commonly exist when deep learning methods are used in real-world scenarios, where the training and test data follow different distributions. They typically stem from factors such as covariate shift in input features, sampling bias in data collection, or concept drift in dynamic systems. To systematically study this phenomenon, existing works often simulate distribution changes. For example, in computer vision, it is common to use images captured under varied conditions as source and target datasets, thereby simulating distribution shifts encountered in real-world scenarios (Glorot et al.,, 2011). There are also works that sample long-tail data, so as to simulate the distribution changes (Jamal et al.,, 2020).
Recently, strategies that simulate distribution changes have been introduced into bioinformatics to better capture realistic scenarios. Lo-Hi (Steshin,, 2023) proposes drug split strategies to simulate lead optimization and hit identification tasks in real drug discovery process. SPECTRA (Ektefaie et al.,, 2024) evaluates model generalizability by plotting performance against decreasing cross-split overlap and reporting the area under this curve. DataSAIL (Joeres et al.,, 2025) models data split as a combinatorial optimization problem and find leakage-reduced data splits based on optimization strategies.
@@ 7 Conclusion and Future Works
In this work, we propose DDI-Ben, a benchmark for emerging DDI prediction in a perspective from drug distribution changes. We propose a distribution change simulation framework that is compatible with different drug split strategies to reflect the distribution change in real-world emerging DDI prediction problems. Through extensive benchmarking evaluation, we find that the distribution changes can cause significant performance drops for existing DDI methods and LLM based methods exhibiting superior robustness compared to other methods. To facilitate future research, we release benchmark datasets that include simulated distribution changes. For future work, designing LLM based methods can potentially alleviate the negative impact of distribution changes in emerging DDI prediction. Domain adaptation methods can also be a potential research direction for practical solutions to deal with distribution changes in real-world emerging DDI prediction.

### v5

@@ 3.3 Customized Distribution Change Surrogate
Split strategy Random Drug frequence Drug usage Scaffold LoHi SPECTRA DataSAIL Cluster
Preserve all data × ×
@@ 5.1 Performance Analysis among Representative Methods
Method Distribution Major Medium Long-tail
@@ 6 Related Works
Simulation of Distribution Changes in Machine Learning. Distribution changes commonly exist when deep learning methods are used in real-world scenarios, where the training and test data follow different distributions. They typically stem from factors such as covariate shift in input features, sampling bias in data collection, or concept drift in dynamic systems. To systematically study this phenomenon, existing works often simulate distribution changes. For example, in computer vision, it is common to use images captured under varied conditions as source and target datasets, thereby simulating distribution shifts encountered in real-world scenarios (Glorot et al.,, 2011). There are also works that sample long-tail data, so as to simulate the distribution changes (Jamal et al.,, 2020). Recently, strategies that simulate distribution changes have been introduced into bioinformatics to better capture realistic scenarios. Lo-Hi (Steshin,, 2023) proposes drug split strategies to simulate lead optimization and hit identification tasks in real drug discovery process. SPECTRA (Ektefaie et al.,, 2024) evaluates model generalizability by plotting performance against decreasing cross-split overlap and reporting the area under this curve. DataSAIL (Joeres et al.,, 2025) models data split as a combinatorial optimization problem and find leakage-reduced data splits based on optimization strategies.
@@ 7 Conclusion and Future Works
In this work, we propose DDI-Ben, a benchmark for DDI prediction in a perspective from drug distribution changes. We propose a distribution change simulation framework that is compatible with different drug split strategies to reflect distribution changes in real-world emerging DDI prediction problems. Through extensive benchmarking evaluation, we find that the distribution changes can cause significant performance drops for existing DDI methods and LLM based methods exhibiting superior robustness. To facilitate future research, we release benchmark datasets that include simulated distribution changes. For future work, designing LLM based methods can potentially alleviate the negative impact of distribution changes in emerging DDI prediction. Domain adaptation methods can also be a potential research direction for practical solutions to deal with distribution changes in real-world emerging DDI prediction.
Acknowledgments
We extend our sincere gratitude to Yang Cao for his assistance in manuscript preparation. We also thank the anonymous reviewers for their insightful comments and constructive suggestions.
Funding
Q. Yao is supported by National Natural Science Foundation of China (under Grant No.92270106) and Beijing Natural Science Foundation (under Grant No.4242039).
Y. Zhang is supported by Guangdong Basic and Applied Basic Research Foundation (Under Grant 2025A1515010304), and Guangzhou Science and Technology Planning Project 2025A03J4491.